jagr2713 (jagr2713):

(143(5) + 67)5 ___________ 2

jagr2713 (jagr2713):

Last one so lets see who is smart :D

pooja195 (pooja195):

143*5=715 +67*5=3910 >:) am i right? ?

jagr2713 (jagr2713):

Nope @pooja195 keep trying though

pooja195 (pooja195):

143*5=715 +67*5=3910/2=1955

OpenStudy (anonymous):

first multiply 143 by 5, then add that answer to 65, then take the answer of all of that and multiply it by 5

pooja195 (pooja195):

The 2 is hidden :o

jagr2713 (jagr2713):

Dang.......... how did you see the 2

jagr2713 (jagr2713):

Sneaky huh

OpenStudy (anonymous):

ah i didnt notice the two

pooja195 (pooja195):

LOL http://prntscr.com/7mz1nm

OpenStudy (anonymous):

quite sneaky

jagr2713 (jagr2713):

Yep almost had you guys MWAHAH

pooja195 (pooja195):

this one is cool :O

jagr2713 (jagr2713):

Lol i thought it would be a little hard and if you guys havent noticed lol I asked this question already :D

pooja195 (pooja195):

o.O wow XD

jagr2713 (jagr2713):

Wow you solved my riddles that took users 15+ to do :D

pooja195 (pooja195):

;)

OpenStudy (usukidoll):

xD! I saw the 2 way at the end XD

OpenStudy (mathstudent55):

(143(5) + 67)5 = (715 + 67)5 =(782)5 =3910

jagr2713 (jagr2713):

you forgot the 2

OpenStudy (mathstudent55):

Since when does a 2 at the end of a line mean divide by 2?

jagr2713 (jagr2713):

Yep its hidden :D

jigglypuff314 (jigglypuff314):

should've done just enough spaces and put a "/2" at the very end instead "__" are kinda like subscripts sometimes ;) ;)

OpenStudy (anonymous):

http://prntscr.com/7n8qsq

jigglypuff314 (jigglypuff314):

That depends on screen size http://prntscr.com/7nipbz @TheЅmartOne